# This Candidate is [MASK] Prompt-based Sentiment Extraction and Reference Letters

### v2

@@ 4.3 Random Forest
In order to explore nonlinear relationships and interactions, I estimate a random forest classifier with the baseline outcome as target and the same predictor variables. [^39] See Athey and Imbens (2019) for details on random forests. The main hyper-parameters are set based on 5-fold cross-validation. [^40] Table A.5 in the appendix presents the hyper-parameter values. As usual, the training of each decision tree is based on a different sample bootstrap and a random subset of the predictors. I apply inverse probability weighting to address the imbalance in the outcome variable. The random forest is trained on the complete applications sub-sample.
One advantage of random forests is that they provide an easy way to assess the predictive performance of the model without the need to set aside a holdout sample. Specifically, it is possible to evaluate each tree in the forest by assessing its predictive performance for the sample elements that were not included in the bootstrap sample used to train the tree. This is referred to as ‘out-of-bag’ (OOB) performance.
@@
Figure 6 presents the relative importance of different predictive features. Specifically, I use permutation importances, which measure the average decrease in the OOB accuracy score when the corresponding feature is shuffled. [^41] Specifically, the feature is re-shuffled and the random forest re-trained 30 times. The average RL sentiment score is the most important predictive feature, with a median decrease in accuracy above 6pp. Average RL length and the sentiment range come second and third, respectively.
I use the trained random forest to obtain partial dependence curves, which simulate the changes in the predicted probability of the outcome for counterfactual values of each feature. Formally, the partial dependence curve for the predictor $X_{S}$ is obtained as follows:
@@
where $X_{C}$ are all other predictors and $\mathcal{X}_{S}$ is the support set of $X_{S}$. It is important to emphasize that the $\mathrm{pd}$ curves try to approximate the conditional expectation function without assuming linearity. They do not provide causal estimates, which can only emerge from an experiment.
In figure 7, I present $\mathrm{pd}$ curves for the main predictors: average RL length, average RL sentiment, and sentiment range. For comparison, I also plot the corresponding simulation based on a linear regression. [^42] The regression in the plot is not exactly the same as the regression in the first column of table 10 because I use inverse probability weights in order to match the random forest fit.
While the random forest confirms the positive associations between RL length, RL sentiment scores and employment outcomes, it suggests that the relationship might not be linear. In particular, the curves seem to have steep jumps after crossing certain thresholds. The random forest also point to the relationship between RL sentiment range and job market success being less pronounced than the regression coefficients suggest.
@@
The plots in figure 9 present similar simulations, this time by assigning counterfactual levels to Avg. RL sentiment and four different discrete predictors: [^43] Similar plots for interactions with Avg. RL length and Sentiment Range are omitted to save space. These and other $\mathrm{pd}$ plots are available in the online appendix (TODO).
@@
4.4 Comparison of Prompt-based Sentiment with Other Approaches
Figure 10: Correlation Matrix
Figure 10 presents Pearson correlations among the different sentiment measures. While almost all correlations are positive, it is noteworthy that the coefficients are clustered by technique. For example, lexicon-based measures have much stronger associations among each other—correlations close to 0.5—than with any of the LLM-based measures. Similarly, the correlation between the two instruction-based measures is 0.82.
I repeated the regression analysis from section 4.1 for each of the alternative sentiment measures. Figure 11 summarizes my findings. Specifically, the box plots in the figure show the p-value distributions for the Avg. RL Sentiment coefficient in equation (3). For each sentiment measure, I estimated 150 different specifications by varying the outcome variable (5 variants), considering alternative control variable sets (5), using either the full or the restricted ‘complete applications sample’ (2), and applying different assumptions to the error term (3).
Somewhat predictably, neither the lexical approach nor the fine-tuned LLMs are able to predict job market outcomes. The dictionaries involved in the former and the fine-tuning process in the latter target applications that are quite distant from RL sentiment extraction.
Regardless of these limitations, the sentiment scores that result from applying ModernBert can contribute to predicting job market outcomes. In most of the regressions, the average RL sentiment comes out significant at the 5% level. [^45] I also trained a random forest using the ModernBert score and found that it was the most important feature in predicting the baseline outcome.
Notes: The box plots depict the p-value distributions for the coefficient corresponding to Avg. RL Sentiment in equation (3). Each sentiment measure is used in 150 alternative specifications: 5 outcome variables, 2 samples, 5 control variable groups, and 3 alternative clusterings of the errors.
Notes: The box plots depict the p-value distributions for the coefficients corresponding to Avg. RL Sentiment and RL Sentiment Range in equation (4). Each sentiment measure is used in 75 alternative specifications estimated on the ‘full applications’ restricted sample: 5 outcome variables, 5 control variable groups, and 3 alternative clusterings of the errors.
Figure 12 repeats the exercise, this time for equation (4). All regressions are estimated on the ‘full application’ restricted sample. Neither ModernBert nor the instruction-based models are able to replicate the findings in table 10.
@@ 5 Gendered Sentiment
Notes: NN=1968. The box plots show the distribution of the net-standout score for female and male candidates.
Are there differences in net-standout scores between male and female candidates? As is clear from figure 13, the answer is a rotund ‘yes’. Similar to the literature, I find that RLs written for male candidates emphasize ‘standout’ traits more, whereas females are more likely to be described using ‘grindstone’ attributes. Moreover, the gender effect is robust to introducing controls for other candidates characteristics, for the PhD rank and region, for the job market year, and for a host of letter writer characteristics including sex, whether the writer is the main adviser, whether the writer is either a top-10 or a top-5 percentile author, and the writer’s title and potential experience. [^48] These regression results are available in the online appendix (TODO).
@@
Using a random forest, I once again explore interactions, this time between the candidate’s sex and other potential predictors of the net-standout score. [^49] Full details on this procedure, including hyper-parameters and performance metrics are available in the online appendix (TODO). The top panel of figure 14 shows the interaction between candidate sex and the rank of their PhD program. The gender gap in net-standout is present throughout the ranking distribution. The table implies that a male candidate from a PhD program ranked 76–100 is likely to get a higher net-standout score than a female in a top-25 program.
@@
In appendix table A.6, I present regression results when equation (3) is augmented to include ‘gendered’ sentiment variables. Specifically, I include the average standout ($\mathrm{P}\left(y=\text{standout}\mid\boldsymbol{x}\right)$) and average grindstone ($\mathrm{P}\left(y=\text{grindstone}\mid\boldsymbol{x}\right)$) scores as predictors of job market outcomes. The coefficients for average letter length and sentiment are almost identical to those in table 8. The coefficients for the average standout score are relatively small and statistically insignificant. The average grindstone score, on the other hand, has a negative and statistically significant association with the probability of success. [^50] The exception is again the outcome that excludes post-docs from the definition of success.
@@ 3.1 Bert Score 4/16E 3
76–100 -0.1286 Job Market Year (omitted)
@@
Four Letters & Avg. Letter Length Yes Yes Yes

### v3

@@ 4.5 Random Forest
Exactly how predictive of success are the contents of RLs? In order to explore this issue, I estimate a random forest classifier with the baseline outcome as target. Besides average RL sentiment and sentiment dispersion, I include all the control variables as predictors. [^43] See Athey and Imbens (2019) for details on random forests. The main hyper-parameters are set based on 5-fold cross-validation. [^44] Table A.5 in the appendix presents the hyper-parameter values. As usual, the training of each decision tree is based on a different sample bootstrap and a random subset of the predictors. I apply inverse probability weighting to address the imbalance in the outcome variable. The random forest is trained on the ‘complete applications’ sub-sample.
Random forests provide an easy way to assess the predictive performance of the model without the need to set aside a holdout sample. Specifically, it is possible to evaluate each tree in the forest by assessing its predictive performance on the sample elements that were not included in the bootstrap sample used to train the tree. This is referred to as ‘out-of-bag’ (OOB) performance.
@@
Figure 9 presents the relative importance of different predictive features. Specifically, I use permutation importances, which measure the average decrease in the OOB accuracy score when the corresponding feature is shuffled. [^45] Specifically, the feature is re-shuffled and the random forest re-trained 30 times. The average RL sentiment score is the most important predictive feature, with a median decrease in accuracy above 6 p.p. Average RL length and the sentiment range come second and third, respectively. Remarkably, other predictor variables usually thought to be very important (the PhD rank, publications, the top-ranked letter writers, having a ‘good job market year’) are not highly predictive of success.
The random forest is also a way to relax the linearity assumption. Specifically, I use the trained forest to obtain partial dependence curves, which simulate the changes in the predicted probability of the outcome for counterfactual values of each feature. Formally, the partial dependence curve for the predictor $X_{S}$ is obtained as follows:
@@
where $X_{C}$ are all other predictors and $\mathcal{X}_{S}$ is the support set of $X_{S}$. It is important to emphasize that the $\mathrm{pd}$ curves try to approximate the conditional expectation function without assuming linearity. They do not provide causal estimates.
In figure 10, I present $\mathrm{pd}$ curves for average RL sentiment and sentiment range. For comparison, I also plot the corresponding simulation based on a linear regression. [^46] The regression in the plot is not exactly the same as the regression in the first column of table 10 because I use inverse probability weights in order to match the random forest fit. While the random forest confirms the positive association between average sentiment scores and employment outcomes, it suggests that the relationship might not be linear. In particular, the curve seems to have steep jumps after crossing certain thresholds. The random forest also points to the relationship between RL sentiment range and job market success being less pronounced than the regression coefficients suggest.
@@
The plots in figure 12 present similar simulations, this time by assigning counterfactual levels to Avg. RL sentiment and four different discrete predictors:
@@ 5 Gendered Sentiment
Notes: NN=1968. The box plots show the distribution of the net-standout score (in p.p.) for female and male candidates. The difference in means is $2.82$ and the HC3 standard error is $0.376$.
Are there differences in net-standout scores between male and female candidates? As is clear from figure 13, the answer is a rotund ‘yes’. Similar to the literature, I find that RLs written for male candidates emphasize ‘standout’ traits more, whereas females are more likely to be described using ‘grindstone’ attributes. Moreover, the gender effect is robust to introducing controls for other candidates characteristics, for the PhD rank and region, for the job market year, and for a host of letter writer characteristics including gender, whether the writer is the main adviser, whether the writer is either a top-10 or a top-5 percentile author, and the writer’s title and potential experience. [^49] These regression results are available in the online appendix (TODO).
@@
Using a random forest, I once again explore interactions, this time between the candidate’s gender and other potential predictors of the net-standout score. [^50] Full details on this procedure, including hyper-parameters and performance metrics are available in the online appendix (TODO). The top panel of figure 14 shows the interaction between candidate gender and the rank of their PhD program. The gender gap in net-standout is present throughout the ranking distribution. The table implies that a male candidate from a PhD program ranked 76–100 is likely to get a higher net-standout score than a female in a top-25 program.
@@
In appendix table A.6, I present regression results when equation (3) is augmented to include ‘gendered’ sentiment variables. Specifically, I include the average standout ($\mathrm{P}\left(y=\text{standout}\mid\boldsymbol{x}\right)$) and average grindstone ($\mathrm{P}\left(y=\text{grindstone}\mid\boldsymbol{x}\right)$) scores as predictors of job market outcomes. The coefficients for average RL sentiment are almost identical to those in table 8. The coefficients for the average standout score are relatively small and statistically insignificant. The average grindstone score, on the other hand, has a negative and statistically significant association with the probability of success. [^51] The exception is again the outcome that excludes post-docs from the definition of success.
@@ 3.1 Bert Score 4/16E 3
76–100 -0.1286 Job Market Year (omitted)#
@@
Four Letters & Avg. RL Length Yes Yes Yes